# Perils of Label Indeterminacy: A Case Study on Prediction of Neurological Recovery After Cardiac Arrest

### v1

@@ Abstract
1 Introduction
@@
AI systems are commonly used for decision support in high-stakes domains, such as healthcare. This typically requires training of supervised AI models. The evaluation of supervised AI models generally assumes access to ground-truth labels against which predictions or decisions can be assessed. If we do not have access to such labels, this can undermine both training and evaluation of AI systems. Unfortunately, ground-truth labels are not always available; they can be missing, incorrect, biased, or uncertain. Research has studied the problem of missing labels [78, 77] as well as the risks of bias and errors contained in labels [37, 24]. In this work, we tackle a related, pervasive, yet understudied problem: there is often an absence of a single, agreed-upon set of ground-truth labels. Instead, labels are often estimated or selected through processes that involve unverifiable assumptions and arbitrary choices—which can have crucial implications. We term this phenomenon label indeterminacy and define it as follows: label indeterminacy refers to situations where practitioners cannot make a principled choice among different options to obtain labels, and as a result are faced with unresolvable multiplicity.
@@
There are multiple conditions that lead to label indeterminacy. When building AI systems to support experts or automate decisions, it is frequently the case that historical observations are used to train them. For instance, a bank may use historical data of loan repayment to train a model for assisting agents in determining who should receive a loan [4]. In this case, outcome labels are only observed conditioned on certain decisions. Concretely, we only observe whether someone can repay a loan if that person was issued a loan in the first place. For rejected applicants, we cannot determine what might have occurred if they had been granted the loan. If we wanted to train an AI model that predicts whether someone will repay a loan, we cannot readily use data from rejected applicants because we do not have ground-truth labels for them [39]. To use such data for training, these instances could be ignored, counterfactual labels could be estimated, and both these choices could be paired with sampling bias correction methods. Choosing among these options, and executing any of them, involves making strong and often unverifiable assumptions [15]. Alternatively, rather than relying on historical data, one may have access to labels that were purposefully collected for model training. For instance, crowdsourcing of labels has been proposed for a variety of tasks, including toxicity detection in online content [27] or for medical relation extraction [18]. In such cases, it is frequently the case that multiple annotations are collected for each instance. As a result, different ways of aggregating annotations may result in different labels. Prior research has alluded to the epistemic and ethical implications of these choices, noting that factors such as annotator identity and knowledge can be weighed to obtain different estimations [27, 44].
@@
In the present work, we study the implications of label indeterminacy for AI-assisted decision-making in high-stakes settings. We do so via a case study in the healthcare domain, focusing on predicting patient recovery after cardiac arrest—a task that has received significant interest in critical care and resuscitation science [55, 31, 38, 47]. In this context, one needs to deal with indeterminate labels for patients who died after withdrawal of life-sustaining therapies. This occurs when, based on clinical circumstances and individual judgments, a healthcare provider believes there is a poor prognosis for recovery, making continued aggressive care unlikely to confer benefit. We model ten different methods of handling indeterminate labels, based on approaches that have been used in the literature or that would be plausible to use. We outline both the potential arguments in favor of their use and the unverifiable assumptions they necessitate. The considered labels include both estimations based on historical decisions and estimations based on an international panel of experts in neurological prognostication. We show that these ten different methods yield models that perform similarly on the subset of cases with known labels (i.e., where recovery is actually observed or patients died despite maximal support), while issuing fundamentally different predictions for patients with indeterminate labels.
@@
Our findings have significant implications. The machine learning literature has focused on proposing variants for counterfactual estimation and label aggregation, showing modeling improvements under certain assumptions. However, as we show in our case study, in the real world these assumptions are often unverifiable or known to be violated. As a result, system designers are forced to make arbitrary choices that have profound implications, raising ethical concerns: a different arbitrary choice can lead to differences in AI recommendations for decisions that have irreversible outcomes. Empirically, we find that in 19.6% of patients in our testing sample with unknown labels, at least one method of handling indeterminate labels predicts no chance of recovery, while another method predicts a moderate (or better) recovery probability of over 10%. When these predictions guide clinical decisions as to whether life-sustaining therapies should be withdrawn, the implications cannot be understated.
@@
Moreover, we show that the perils of label indeterminacy are exacerbated by the fact that the different approaches may yield comparable performance when assessed on patients with known labels, i.e., the subset of patients where performance can be directly assessed. This means that these differences may go undetected, and that it may not be possible to discard approaches based on poor performance on this subset. As we argue in this work, this results in a form of multiplicity that undermines the viability of these models for use in high-stakes settings. Thus, efforts to improve modeling of missing labels do not address a fundamental problem: the indeterminate nature of these labels. As a path forward, we advocate for novel sociotechnical designs that shift the objective of prediction to other constructs that do not suffer from indeterminacy but are still useful for humans to reach high-quality decisions.
2 Background and Related Work
@@
True labels can be missing for different reasons. One reason of particular importance to our case study is the so-called selective labels problem [39, 72]. Under selective labels, we only have a partial labeling of instances, and whether we observe true labels depends on historical decisions—such as whether or not life-sustaining therapies are withdrawn [14]. When ground-truth labels are unavailable for training, they must be directly or indirectly estimated [50]. This is a problem that has received substantial attention in the statistics and machine learning literature. Prior work has suggested different techniques for estimating counterfactuals, such as nearest-neighbor imputation [10], which require assumptions regarding the relevance of a given distance metric. Other works treat missing labels as a sampling bias problem and aim to correct for it. This typically requires assuming that every case has a non-negligible probability of having a label observed [15] as well as that no unobservables affect the likelihood of observing a label [39], which are often unrealistic. Another approach, which is ubiquitous in biomedical research [14], is to assume that historical decisions that led to not observing a true label were correct, and to impute the labels accordingly. Training AI models on these labels can, however, result in predictions that mimic potentially erroneous historical decisions and lead to harmful self-fulfilling prophecies [14, 21]. Taken together, prior work has proposed many different approaches to impute missing labels and correct for sampling bias, albeit at the cost of making strong assumptions. Through our case study, we show that choosing among the many available options often becomes an arbitrary choice, because each makes assumptions that are unverifiable or violated in practice. As a result, claims of improved modeling abilities do not address the fundamental challenges encountered in high-stakes decision-making.
@@
While in some cases, labels are missing for some instances, in others, there may be multiple labels for each instance. Often, multiple estimates are collected for a single instance, with the goal of obtaining higher-quality labels [35, 65]. A similar challenge emerges when convening expert panels, which are often considered the gold standard in generating high-quality labels [51, 75]. This raises the question of how to aggregate multiple assessments for a single instance, and how to deal with expert disagreement [69, 27, 70]. For that purpose, different methods have been proposed, ranging from majority voting [59] to more involved techniques like jury learning [27]. Crucially, if expert assessments correspond to fundamentally different worldviews, prior work has argued that reconciling them in a single aggregate value appears nonsensical [44]. In this work, we have the unique opportunity to make use of a dataset containing estimates from a panel of 38 world-leading experts in neurological prognostication to assess patient cases with respect to their recovery potential. This allows us to study the real-world implications of different choices of aggregating this information in high-stakes decision-making when faced with label indeterminacy, compare the option of learning solely from historical data vs. integrating panel assessments, and highlight the ethical challenges these choices pose.
@@
Prior work has noted the fact that available labels may suffer from construct validity issues [36, 56], i.e., a gap between the construct of interest and what is captured in proxies used as training labels [16]. Some work has treated this as a statistical modeling challenge, aiming to “bridge the construct gap” [16]. Other work has noted that treating it as part of the problem formulation stage, and empirically assessing the implications of different alternatives, may guide the correct choice [53]. Label indeterminacy may or may not arise as a result of construct validity issues. For instance, the outcome observed may correspond to the construct of interest, but it may not always be available. Furthermore, we place emphasis on the indeterminate nature of this problem—meaning, it cannot be determined. We empirically show that making a principled choice among options may be impossible, forcing designers to grapple with the ethical implications of choosing arbitrarily. Through this conceptual difference, we are able to elucidate a crucial problem in the design of AI systems for decision support, and propose a path forward through a sociotechnical lens.
@@
Multiplicity is the phenomenon that different (potentially arbitrary) design choices in the development of AI systems can lead to different outcomes. Various types of multiplicity have been proposed and empirically studied in prior work. Predictive multiplicity refers to the situation where a prediction problem admits competing models with conflicting predictions [46, 71, 5]. In these cases, a single dataset can support multiple solutions that are equivalent on the basis of relevant established performance metrics. Brunet et al. [7] highlight implications of multiplicity for explanations, and Coston et al. [11] for fairness. All these notions of multiplicity have in common that they assume access to a single set of ground-truth labels. This is an important conceptual difference to our work, where we have no grounds for determining what is correct for at least a portion of instances.
@@
The degrees of freedom for system designers afforded by multiplicity bring about the possibility to prioritize other values in the model selection process besides predictive accuracy, but may also lead to problems around justifiability and cause loopholes where system designers can mask intentional discrimination [5]. In this work, we show that label indeterminacy can lead to a form of multiplicity when multiple possible avenues to construct labels are equally justifiable. We empirically show that this may result in settings where AI predictions differ drastically for instances with unknown labels, while maintaining comparable performance for those with known labels. This has critical implications for decision subjects and poses challenges for current practices in AI-assisted decision-making that have received limited attention in previous research.
3 Label Indeterminacy
@@
AI systems for decision automation or support often rely on supervised machine learning algorithms, which require the availability of a “ground-truth” label $Y$. In many applications, however, such a label is often the result of a process spanning problem formulation [56] and data engineering [54]. Obscured in the end result is the fact that many other possible labels could have emerged from that process, and arbitrary choices in formulation and estimation led to an implicit selection among them. We introduce the term label indeterminacy to refer to situations where at least a portion of the ground-truth labels are unknowable and different ways of estimating them involve unverifiable assumptions or arbitrary choices, rendering it impossible to make a principled selection among them.
@@
Intertwined in this notion are two important components: one that relates to the properties of the world itself and one that relates to our statistical estimation capabilities. The unknowable nature of the labels is a notion grounded on philosophy literature, which posits that indeterminacy is a property of the world itself rather than what we know about the world [60]. Hence, it is not something we can aim to resolve by acquiring new knowledge or information. This refers to the label for a specific, individual instance. For instance, if life-sustaining therapies are withdrawn from a critically ill patient and as a result the patient dies, it can never be known whether that patient would have recovered had life-sustaining therapies been continued. This unknowable attribute is present across domains, and it is an essential property of settings in which the question relates to counterfactuals: we can never know what might have occurred if events had unfolded differently. In the context of machine learning, however, the quantity of interest is not the estimation for an instance in the past, but rather the prediction for an instance in the future. In the example of estimating recovery, a label indicating what would have happened had we continued life-sustaining therapies is a means to an end: providing better prognosis for patients who have not yet been treated. Consequently, we are interested in label estimation that yields reliable and accurate predictions for new patients. The problem, however, emerges when multiple such labels exist, and it is impossible to make a principled choice among them, because they each make assumptions that are either themselves unknowable or known to be false. In such cases, the quantity of interest is indeterminate.
@@
Formally, let $Y_{i}$ be a ground-truth label for instance $i\in I$, with $I=\{1,...,m\}$, $m\in\mathbb{N}$. $Y_{i}$ can be known or unknown at the time of dataset construction. If $Y_{i}$ is unknown, in some cases one can collect more empirical evidence to arrive at the true label; for instance, by means of human annotators [17]. In other cases, $Y_{i}$ cannot be known—which is the situation of interest in this work. For these instances, there may be a set of $\ell$ plausible labels $\{Y_{i}^{1},Y_{i}^{2},...,Y_{i}^{\ell}\}$, and knowing which among them is more likely to be correct may be impossible. We refer to the subset of instances $W\subseteq I$ with unknowable labels as label indeterminacy set.
@@
Indeterminacy poses a fundamental problem for machine learning when the consequences of selecting one among the available labels deem it unethical to let this be an arbitrary choice. Next, we empirically study label indeterminacy in a high-stakes setting, and derive implications for the development and deployment of AI for decision support.
4 Empirical Study

### v2

@@ Abstract.
Label indeterminacy, multiplicity, AI-assisted decision-making, healthcare, case study
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732070isbn: 979-8-4007-1482-5/2025/06ccs: Computing methodologies Machine learningccs: Information systems Decision support systemsccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
@@
AI systems are commonly used for decision support in high-stakes domains, such as healthcare. This typically requires training of supervised AI models. The evaluation of supervised AI models generally assumes access to ground-truth labels against which predictions or decisions can be assessed. If we do not have access to such labels, this can undermine both training and evaluation of AI systems. Unfortunately, ground-truth labels are not always available; they can be missing, incorrect, biased, or uncertain. Research has studied the problem of missing labels (Zschech et al., 2019; Zhu and Goldberg, 2022) as well as the risks of bias and errors contained in labels (Jiang and Nachum, 2020; Fogliato et al., 2020). In this work, we tackle a related, pervasive, yet understudied problem: there is often an absence of a single, agreed-upon set of ground-truth labels. Instead, labels are often estimated or selected through processes that involve unverifiable assumptions and arbitrary choices—which can have crucial implications. We term this phenomenon label indeterminacy and define it as follows: label indeterminacy refers to situations where practitioners cannot make a principled choice among different options to obtain labels, and as a result are faced with unresolvable multiplicity.
@@
There are multiple conditions that lead to label indeterminacy. When building AI systems to support experts or automate decisions, it is frequently the case that historical observations are used to train them. For instance, a bank may use historical data of loan repayment to train a model for assisting agents in determining who should receive a loan (Bao et al., 2019). In this case, outcome labels are only observed conditioned on certain decisions. Concretely, we only observe whether someone can repay a loan if that person was issued a loan in the first place. For rejected applicants, we cannot determine what might have occurred if they had been granted the loan. If we wanted to train an AI model that predicts whether someone will repay a loan, we cannot readily use data from rejected applicants because we do not have ground-truth labels for them (Lakkaraju et al., 2017). To use such data for training, these instances could be ignored, counterfactual labels could be estimated, and both these choices could be paired with sampling bias correction methods. However, choosing among these options, and executing any of them, involves making strong and often unverifiable assumptions (De-Arteaga et al., 2018). Alternatively, rather than relying on historical data, one may have access to labels that were purposefully collected for model training. For instance, crowdsourcing of labels has been proposed for a variety of tasks, including toxicity detection in online content (Gordon et al., 2022) or for medical relation extraction (Dumitrache et al., 2018). In such cases, it is frequently the case that multiple annotations are collected for each instance. As a result, different ways of aggregating these annotations may result in different labels.
The machine learning literature has typically treated different sources of label indeterminacy as belonging to separate streams of research (e.g., missing labels vs. aggregating multiple labels). However, as we demonstrate in our work, these challenges can be intertwined in practice, since the arbitrary choices practitioners must make can lead to different problem formulations. For instance, in settings where outcomes are selectively observed, one may choose to constrain oneself to observed labels and treat the problem as sampling bias, or collect expert assessments and treat it as a label aggregation problem.
@@
In the present work, we study the implications of label indeterminacy for AI-assisted decision-making in high-stakes settings. We do so via a case study in the healthcare domain, focusing on predicting patient recovery after cardiac arrest—a task that has received significant interest in critical care and resuscitation science (Park et al., 2019; Hessulf et al., 2023; Kwon et al., 2019; Mayampurath et al., 2022). In this context, one needs to deal with indeterminate labels for patients who died after withdrawal of life-sustaining therapies. This occurs when, based on clinical circumstances and individual judgments, a healthcare provider believes there is a poor prognosis for recovery, making continued aggressive care unlikely to confer benefit. We model ten different methods for handling indeterminate labels, based on approaches that have been used in the literature or that would be plausible to use. We outline both the potential arguments in favor of their use and the unverifiable assumptions they necessitate. The considered labels include both estimations based on historical decisions and estimations based on an international panel of experts in neurological prognostication. We show that these ten different methods yield models that perform similarly on the subset of cases with known labels (i.e., where recovery is actually observed or patients died despite maximal support), while issuing fundamentally different predictions for patients with indeterminate labels.
@@
Our findings have significant implications. The machine learning literature has focused on proposing variants for counterfactual estimation and label aggregation, showing modeling improvements under certain assumptions. However, as we show in our case study, in the real world these assumptions are often unverifiable or known to be violated. As a result, system designers are forced to make arbitrary choices that have profound implications, raising ethical concerns: a different arbitrary choice can lead to differences in AI recommendations for decisions that have irreversible outcomes. Empirically, we find that in 19.6% of patients in our testing sample with unknown labels, at least one method of handling indeterminate labels predicts no chance of recovery, while another method predicts at least a moderate recovery probability of over 10%. When these predictions guide clinical decisions regarding the continuation or withdrawal of life-sustaining therapies, the consequences are profound.
@@
Moreover, we show that the perils of label indeterminacy are exacerbated by the fact that the different approaches may yield comparable performance when evaluated on patients with known labels, i.e., the subset of patients where performance can be directly assessed. This means that these differences in predictions may go undetected, and that it may not be possible to discard approaches based on poor performance on the subset of instances with known labels. As we argue in this work, this results in a form of multiplicity that undermines the viability of these models for use in high-stakes settings. Thus, efforts to improve modeling of missing labels do not address a fundamental problem: the indeterminate nature of these labels. As a path forward, we advocate for novel sociotechnical designs that shift the objective of prediction to other constructs that do not suffer from indeterminacy but are still useful for humans to reach high-quality decisions.
2. Background and Related Work
@@
True labels can be missing for different reasons. One reason of particular importance to our case study is the so-called selective labels problem (Lakkaraju et al., 2017; Wei, 2021). Under selective labels, we only have a partial labeling of instances, and whether we observe true labels depends on historical decisions—such as whether or not life-sustaining therapies are withdrawn (De-Arteaga and Elmer, 2023). Other examples of missing labels include sample dropout in medical trials (Hogan et al., 2004; Little and Yau, 1996) or so-called positive and unlabeled data (PU), where only certain labels are recorded (Bekker and Davis, 2020; Wang et al., 2025; Perini et al., 2023).
When ground-truth labels are unavailable for training, they must be directly or indirectly estimated (Muller et al., 2021). This is a problem that has received substantial attention in the statistics and machine learning literature. Prior work has suggested different techniques for estimating counterfactuals, such as nearest-neighbor imputation (Chen and Shao, 2000), which require assumptions regarding the relevance of a given distance metric. Other works treat missing labels as a sampling bias problem and aim to correct for it. This typically requires assuming that every case has a non-negligible probability of having a label observed (De-Arteaga et al., 2018) as well as that no unobservables affect the likelihood of observing a label (Lakkaraju et al., 2017), which are often unrealistic. Another approach, which is ubiquitous in biomedical research (De-Arteaga and Elmer, 2023), is to assume that historical decisions that led to not observing a true label were correct, and to impute the labels accordingly. Training AI models on these labels can, however, result in predictions that mimic potentially erroneous historical decisions and lead to harmful self-fulfilling prophecies (De-Arteaga and Elmer, 2023; Elmer et al., 2023).
Taken together, prior work has proposed many different approaches to impute missing labels and correct for sampling bias, albeit at the cost of making strong assumptions. Through our case study, we show that choosing among the many available options often becomes an arbitrary choice, because each makes assumptions that are unverifiable or violated in practice. As a result, claims of improved modeling abilities do not address the fundamental challenges encountered in high-stakes decision-making.
@@
While in some cases labels are missing for some instances, in others there may be multiple labels for each instance. Often, multiple estimates are collected for a single instance, with the goal of obtaining higher-quality labels (Ipeirotis et al., 2014; Sheng et al., 2008). A similar challenge emerges when convening expert panels, which are often considered the gold standard in generating high-quality labels (Nguyen et al., 2015; Willemink et al., 2020). This raises the question of how to aggregate multiple assessments for a single instance, and how to deal with expert disagreement (Uma et al., 2021; Gordon et al., 2022; Wallace et al., 2022). For that purpose, different methods have been proposed, ranging from majority voting (Quoc Viet Hung et al., 2013) to more involved techniques like jury learning (Gordon et al., 2022). Crucially, if expert assessments correspond to fundamentally different worldviews, prior work has argued that reconciling them in a single aggregate value appears nonsensical (Manski et al., 2016). In this work, we have the unique opportunity to make use of a dataset containing estimates from a panel of 38 world-leading experts in neurological prognostication to assess patient cases with respect to their recovery potential. This allows us to study the real-world implications of different choices of aggregating this information in high-stakes decision-making when faced with label indeterminacy, compare the option of learning solely from historical data vs. integrating panel assessments, and highlight the ethical challenges these choices pose.
@@
Prior work has noted the fact that available labels may suffer from construct validity issues (Jacobs and Wallach, 2021; Passi and Barocas, 2019), i.e., a gap between the construct of interest and what is captured in proxies used as training labels (De-Arteaga et al., 2021). Some work has treated this as a statistical modeling challenge, aiming to “bridge the construct gap” (De-Arteaga et al., 2021). Other work has noted that treating it as part of the problem formulation stage, and empirically assessing the implications of different alternatives, may guide the correct choice (Obermeyer et al., 2019). Label indeterminacy may or may not arise as a result of construct validity issues. For instance, the outcome observed may correspond to the construct of interest, but it may not always be available. Furthermore, we place emphasis on the indeterminate nature of this problem—meaning, it cannot be determined. We empirically show that making a principled choice among options may be impossible, forcing designers to grapple with the ethical implications of choosing arbitrarily. Through this conceptual difference, we are able to elucidate a crucial problem in the design of AI systems for decision support, and propose a path forward through a sociotechnical lens.
@@
Multiplicity is the phenomenon that different (potentially arbitrary) design choices in the development of AI systems can lead to different outcomes. Various types of multiplicity have been proposed and empirically studied in prior work. Predictive multiplicity refers to the situation where a prediction problem admits competing models with conflicting predictions (Marx et al., 2020; Watson-Daniels et al., 2023; Black et al., 2022). In these cases, a single dataset can support multiple solutions that are equivalent on the basis of relevant established performance metrics. Brunet et al. (2022) highlight implications of multiplicity for explanations, and Coston et al. (2021) for fairness. All these notions of multiplicity have in common that they assume access to a single set of ground-truth labels. This is an important conceptual difference to our work, where we have no grounds for determining what is correct for at least a portion of instances.
@@
The degrees of freedom for system designers afforded by multiplicity bring about the possibility to prioritize other values in the model selection process besides predictive accuracy, but may also lead to problems around justifiability and cause loopholes where system designers can mask intentional discrimination (Black et al., 2022). In this work, we show that label indeterminacy can lead to a form of multiplicity when multiple possible avenues to construct labels are equally justifiable. We empirically show that this may result in settings where AI predictions differ drastically for instances with unknown labels, while maintaining comparable performance for those with known labels. This has critical implications for decision subjects and poses challenges for current practices in AI-assisted decision-making that have received limited attention in previous research.
3. Label Indeterminacy
@@
AI systems for decision automation or support often rely on supervised machine learning algorithms, which require the availability of a “ground-truth” label $Y$. In many applications, however, such a label is often the result of a process spanning problem formulation (Passi and Barocas, 2019) and data engineering (Orr and Crawford, 2024). Obscured in the end result is the fact that many other possible labels could have emerged from that process, and arbitrary choices in formulation and estimation led to an implicit selection among them. We introduce the term label indeterminacy to refer to situations where at least a portion of the ground-truth labels are unknowable and different ways of estimating them involve unverifiable assumptions or arbitrary choices, rendering it impossible to make a principled selection among them.
@@
Intertwined in this notion are two important components: one that relates to the properties of the world itself and one that relates to our statistical estimation capabilities. The unknowable nature of the labels is a notion grounded in philosophy literature, which posits that indeterminacy is a property of the world itself rather than what we know about the world (Richardson, 2024). Hence, it is not something we can aim to resolve by acquiring new knowledge or information. This refers to the label for a specific, individual instance. For example, if life-sustaining therapies are withdrawn from a critically ill patient and as a result the patient dies, it can never be known whether that patient would have recovered had life-sustaining therapies been continued. This unknowable attribute is present across domains, and it is an essential property of settings in which the question relates to counterfactuals: we can never know what might have occurred if events had unfolded differently. In the context of machine learning, however, the quantity of interest is not the estimation for an instance in the past, but rather the prediction for an instance in the future. In the example of estimating recovery, a label indicating what would have happened had we continued life-sustaining therapies is a means to an end: providing better prognosis for patients who have not yet been treated. Consequently, we are interested in label estimation that yields reliable and accurate predictions for new patients. The problem, however, emerges when multiple such labels exist, and it is impossible to make a principled choice among them, because they each make assumptions that are either themselves unknowable or known to be false. In such cases, the quantity of interest is indeterminate.
@@
Formally, let $Y_{i}$ be a ground-truth label for instance $i\in I$, with $I=\{1,...,m\}$, $m\in\mathbb{N}$. $Y_{i}$ can be known or unknown at the time of dataset construction. If $Y_{i}$ is unknown, in some cases one can collect more empirical evidence to arrive at the true label; for instance, by means of human annotators (Deng et al., 2014). In other cases, $Y_{i}$ cannot be known—which is the situation of interest in this work. For these instances, there may be a set of $\ell$ plausible labels $\{Y_{i}^{1},Y_{i}^{2},...,Y_{i}^{\ell}\}$, and knowing which among them is more likely to be correct may be impossible. We refer to the subset of instances $W\subseteq I$ with unknowable labels as label indeterminacy set.
@@
Indeterminacy poses a fundamental problem for machine learning when the consequences of selecting one among the available labels deem it unethical to let this be an arbitrary choice. Next, we empirically study label indeterminacy in a high-stakes setting, and derive implications for the development and deployment of AI systems for decision support.
4. Empirical Study
Figure 1. Obtaining labels depends on whether patients are withdrawn from life-sustaining therapies (WLST) or not. In the case of WLST, patients die, and we do not know whether they could have recovered had they not been withdrawn from life-sustaining therapies. For WLST cases, labels are unknown.
[…]
An illustration of label indeterminacy in the context of the case study.